### v1

# Deviations from random matrix entanglement statistics for kicked quantum chaotic spin-$1/2$ chains

Abstract
It is commonly expected that for quantum chaotic many body systems the statistical properties approach those of random matrices when increasing the system size. We demonstrate for various kicked spin-$1/2$ chain models that the average eigenstate entanglement indeed approaches the random matrix result. However, the distribution of the eigenstate entanglement differs significantly. While for autonomous systems such deviations are expected, they are surprising for the more scrambling kicked systems. We attribute the origin of the deviations to the local two-dimensional Hilbert spaces. This is also supported by similar deviations occurring in a local random matrix model with global diagonal coupling.
The characterization of statistical properties of complex quantum systems has become an important tool in many areas of physics. Such a statistical description has started with complex nuclei, which can be described by results from random matrix theory (RMT) GuhMueWei1998 ; Meh2004 . Even single-particle systems follow RMT if the corresponding classical dynamics is chaotic, as stated by the famous Bohigas-Giannoni-Schmit conjecture BohGiaSch1984 . Based on a semiclassical trace formula description Gut1990 , this can be explained for the spectral form factor using specific properties of periodic orbits Ber1985 ; SieRic2001 ; Sie2002 ; MueHeuBraHaaAlt2004 ; MueHeuBraHaaAlt2005 ; MueHeuAltBraHaa2009 . More recently, it has become common to call a many-body system quantum chaotic if its level spacing distribution follows the RMT prediction, see e.g. Refs. JenSha1985 ; HsuAng1993 ; MonPoiBelSir1993 ; JacShe1997 ; AviRicBer2002 ; San2004 ; PonPapHuvAba2015 ; KheLazMoeSon2016 ; AkiWalGutGuh2016 ; BorLueSchKnaBlo2017 ; LuiBar2017b ; KosLjuPro2018 . A general expectation is that for a quantum chaotic many body system, also other statistics, like eigenvector distributions or entanglement follow the RMT results Izr1990 ; ZelBroFraHor1996 ; AleKafPolRig2016 .
For autonomous many-body systems states at the edge of the spectrum are well known to differ from RMT EisCraPle2010 and in particular show area law entanglement Sre1993 . However it has recently been found, that even the midspectrum eigenstates do not fulfil the RMT expectations. Deviations from the RMT results are found at the level of eigenvector distributions, such as the components of the eigenvectors SrdProSot2021 and the fractal dimensions BaeHaqKha2019 ; PauCarRodBuc2021 , but also for the entanglement entropy VidHacBiaRig2017 ; VidRig2017 ; BeuBaeMoeHaq2018 ; HaqClaKha2022 . Some of these deviations can be understood analytically Hua2019 ; Hua2021 . An intuitive understanding for the deviations is based on the orthogonality of the eigenstates which implies that the non-RMT states at the edge of the spectrum impose for local Hamiltonians an additional structure on the midspectrum eigenstates HaqClaKha2022 .
The situation is much less understood for Floquet many-body systems such as kicked systems. These systems are more scrambling in the sense that the time evolution operator is less local than in case of autonomous systems. Their dynamics is described by an effective Hamiltonian which have interactions in all possible ranges and orders HerKieBae2023 . Additionally, there is no energy dependence of the statistical properties of eigenstates. Therefore, the theoretical understanding of the autonomous case does not carry over to the kicked situation. Recently it has been shown that for kicked quantum chaotic many-body systems spectral properties such as the spectral form factor are excellently described by RMT at least in the thermodynamic limit KosLjuPro2018 ; BerKosPro2018 ; ChaLucCha2018b ; ChaLucCha2018a ; FriChaDeCha2019 ; FlaBerPro2020 ; RoyPro2020 ; BerKosPro2021 ; GarCha2021 ; ChaDeCha2021 ; FriKie2024 . Also for some average statistical properties of eigenstates agreement with RMT predictions was shown AleRig2014 ; SieLewScaZak2023 .
However, a detailed understanding of the entanglement behavior for quantum chaotic kicked spin chains is still lacking. In particular, will kicked spin chains approach the expected random matrix predictions in the thermodynamic limit of large dimension?
In this paper we establish for the prototypical class of quantum chaotic kicked spin-chains that the average entanglement entropy and the standard deviation approach RMT predictions. However, surprisingly, the distribution of the von Neumann entropy does not approach the distribution for the corresponding RMT ensemble. The distributions stay close but distinct, even for large system sizes, which strongly suggests that this persists also in the thermodynamic limit. This effect also occurs for a simple random matrix model with global coupling. Thus the deviations are not caused by the short range interactions in the kicked spin chains, but by the two-dimensional local Hilbert spaces.
Quantum chaotic many body systems – Many body systems are commonly called quantum chaotic, if their level spacing distribution follows the predictions from RMT JenSha1985 ; HsuAng1993 ; MonPoiBelSir1993 ; JacShe1997 ; AviRicBer2002 ; San2004 ; PonPapHuvAba2015 ; KheLazMoeSon2016 ; AkiWalGutGuh2016 ; BorLueSchKnaBlo2017 ; LuiBar2017b ; KosLjuPro2018 . In the following we restrict ourselves to predictions for the circular orthogonal ensemble (COE), which is the appropriate RMT ensemble for systems with time reversal invariance, such as the kicked spin chains considered for illustrations in this paper.
For a unitary time evolution operator $U$ acting on an $N$ dimensional Hilbert space, the eigenvalue equation reads
$U|\psi_{n}\rangle=\text{e}^{\text{i}\varphi_{n}}|\psi_{n}\rangle\;,\quad\text{with}\quad n=1,2,\dots,N\;.$ (1)
The eigenstates $|\psi_{n}\rangle$ are assumed to be normalized and the eigenphases fulfill $\varphi_{n}\in[-\pi,\pi)$. Furthermore, we choose the phases $\{\varphi_{1},\varphi_{2},\dots,\varphi_{N}\}$ to be ordered increasingly. The (consecutive) level spacing distribution is the distribution of the spacings $s_{n}=\frac{N}{2\pi}\left(\varphi_{n+1}-\varphi_{n}\right)$, with $\varphi_{N+1}:=\varphi_{1}+2\pi$. Here the pre-factor provides the unfolding, leading to a unit mean spacing. In the limit $N\rightarrow\infty$ the COE result of RMT is well-described by the Wigner distribution Wig1967
$P_{\text{COE}}(s)\approx\frac{\pi}{2}s\exp\left(-\frac{\pi}{4}s^{2}\right)\;.$ (2)
It is expected that for quantum chaotic many-body systems the eigenstate entanglement follows the results predicted from RMT Lub1978 ; Pag1993 ; Sen1996 ; KumPan2011 . For a bipartite system the eigenstate entanglement can be quantified using the von Neumann entropy. To obtain this bipartite structure one can split the $N$–dimensional system into two subsystems of dimension $N_{1}$ and $N_{2}$ with $N=N_{1}N_{2}$. In the following, only the case of $N_{1}=N_{2}=\sqrt{N}$ will be discussed for simplicity. The von Neumann entropy for a state $|\psi\rangle$ is defined by
$S_{1}=-\text{tr}(\rho_{1}\ln(\rho_{1}))\;,$ (3)
with $\rho_{1}=\text{tr}_{2}(\rho)$ being the reduced density matrix of subsystem 1, resulting from tracing out the second subsystem from the density matrix $\rho=|\psi\rangle\langle\psi|$. Unentangled states can be written as product states $|\psi\rangle=|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ and have zero entropy, while maximally entangled states have $S_{1}^{\text{max}}=\ln\left(N_{1}\right)$. The average entropy of random states from the COE is slightly reduced. For $1\ll N_{1}\leq N_{2}$ this entropy is given by Pag1993 ,
$S_{1}^{\text{COE,}N_{1}\gg 1}=\ln\left(N_{1}\right)-\frac{N_{1}}{2N_{2}}\;.$ (4)
Note that in the following we have to use the exact formula given in Ref. (KumPan2011, , Eq. (48)), which is valid without any restrictions for $N_{1}$ and $N_{2}$.
Kicked field Ising model – As prototypical system to study entanglement in many body systems, we consider the kicked field Ising model (KFIM) which is given by the time evolution operator Pro2002 ; Pro2007 ; BerKosPro2018 ; BerKosPro2019
$U_{\text{KFIM}}=$ $\text{e}^{-\text{i}H^{z}}\text{e}^{-\text{i}H^{x}}\;,\quad\text{with}$ (5)
$H^{z}=$ $J\sum_{j=1}^{L}\sigma_{j}^{z}\sigma_{j+1}^{z}+\sum_{j=1}^{L}h_{j}\sigma_{j}^{z}$ (6)
$H^{x}=$ $b\sum_{j=1}^{L}\sigma^{x}_{j}\;.$ (7)
Here $\sigma^{x}_{j}$ and $\sigma^{z}_{j}$ are the standard Pauli spin-matrices acting on the $j$th spin and $L$ is the number of spins in the chain so that the dimension of the Hilbert space is $N=2^{L}$. The first part $H^{z}$, Eq. (6), contains the nearest neighbor coupling in the $z$ component with strength $J$ and a position dependent longitudinal field $\{h_{j}\}$. The second part, Eq. (7), represents a transverse magnetic field with kicking strength $b$ which is periodically turned on and off. This form is equivalent to a free evolution with $H^{z}$ (6) and periodic kick described by $H^{x}$ (7). The longitudinal field components $\{h_{j}\}$ are chosen randomly from a normal distribution with mean $\overline{h}$ and standard deviation $\sigma_{h}$. We consider the self dual KFIM ($J=b=\pi/4$), which is time reversal invariant and obeys another antiunitary symmetry BraWalAkiGutGuh2020 ; FlaBerPro2020 . Therefore, the appropriate RMT ensemble is the T+ CRE (BD I in Cartan’s classification Due2004 ). However, for the level spacing distribution and the distribution of the von Neumann entropy, we see no numerical differences between COE and T+ CRE. Thus, for simplicity, we use COE as reference. For numerical computations  [^1] For the three chain models we used $J=\pi/4$, $\overline{h}=0.6$, $\sigma_{h}=\pi/4$ and $b=\pi/4$ respectively $g=\pi/4$. For large dimensions we use the POLFED algorithm SieLewZak2020 ; Lui2021 ; SieLewScaZak2023 with number of eigenvectors calculated $n_{\text{ev}}=\min(2^{L}/10,1000)$, dimension of Krylov space $n_{\text{cv}}=2n_{\text{ev}}$, a pylynomial degree of $k=0.95\cdot 2^{L+1}/n_{\text{cv}}$ and a target phase $\phi_{\text{tgt}}=\pi/2$. POLFED is used for the KFIM and the random matrix model if $L>14$, for $C_{\text{NN}}$ chain and $C_{\text{NNNN}}$ chain if $L>8$. As number of realizations we use for $L=6,8,\dots,20$ for the KFIM $300,250,100,30,20,25,15,4$, for the $C_{\text{NN}}$ chain and the $C_{\text{NNNN}}$ chain $300,250,500,100,30,25,15,4$ and for the random matrix model $500,300,100,50,30,25,15,4$. For the COE we use $25600,76800,204800,204800,491520,327680,20000,10000$ states. we use the polynomially filtered exact diagonalization (POLFED) algorithm SieLewZak2020 ; Lui2021 ; SieLewScaZak2023 for long chains.
Figure 1: (a) Level spacing distribution for the KFIM with $L=16$ (gray) and $L=20$ (black) and analytical predictions (2) (red dashed line). (b) Distribution of von Neumann entropy for the KFIM with $L=16$ (gray histogram) with average value shown as dotted line. Numerical results for random states (red histogram) and analytical prediction for COE (red dashed line). (c) similar to (b) for $L=20$, where data of KFIM is shown in black.
Figure 1(a) shows the level spacing distribution for the KFIM with $L=16$ (gray) compared to the COE prediction (2) (red). We see that the distribution for the KFIM follows the prediction for COE very well. Thus, the KFIM for $L=16$ qualifies as quantum chaotic for the chosen parameters. Consequently, we would expect the von Neumann entropy of the eigenstates to show COE behavior.
Figure 1(b) shows the distribution of the eigenstate entanglement for the KFIM with $L=16$ (gray) as well as the averaged von Neumann entropy as a dotted line. The predicted average value for COE states of corresponding dimension (KumPan2011, , Eq. (48)) is shown as a red dashed line. In lack of an analytical prediction for the distribution of the von Neumann entropy of random states, we compare the distribution for the KFIM with numerical results for COE states. The average von Neumann entropy for the KFIM is close to the COE prediction and also the distributions are close. However, there are clear differences between the results for the KFIM and the COE. A possible explanation for these deviations could be that the system size is not large enough, since we would expect the COE results to hold in the limit of large systems sizes. Thus, we would expect that with increasing $L$ the deviations decreases and the KFIM entropy approaches the COE result.
Therefore, we consider $L=20$, which also qualifies as quantum chaotic according to the level spacing distribution, see Fig. 1(a) black histogram. Figure 1(c) shows the average value and the distribution of the entropy for the KFIM with $L=20$ compared to COE results of appropriate size. One sees, that the average values are close to each other and that the distributions are also close. However, compared to the results for $L=16$ in Fig. 1(b) it is not clear whether the agreement between the KFIM and the COE distribution has improved, i.e. whether the distributions approach each other. Note that we have adjusted the $S_{1}$-axis of the plots (b) and (c) in Fig. 1 such that the distributions of the different lengths can be compared visually. Thus the axis in (c) has a much smaller $S_{1}$-range than in (b). This means that the absolute distance between the average von Neumann entropy of the KFIM and the predicted result for COE (KumPan2011, , Eq. (48)), actually decreases as $L$ increases. However, it is not possible to tell from these plots whether the distributions approach each other or not.
Figure 2: Characterization of the distribution of the von Neumann entropy in depence on the chain length $L$ for different models. (a) Absolute difference between average values and the COE prediction, Eq. (8). (b) Absolute difference between standard deviations the numerically computed COE value, Eq. (9). (c) Ratio of the absolute difference between the average values and the standard deviation, Eq. (10). As models KFIM (black $+$), $C_{\text{NN}}$ chain (green crosses), $C_{\text{NNNN}}$ chain (blue circles), and random matrix model (yellow squares).
Figure 3: Visualization of different values of ratio $\mathcal{R}$ (10). As an example, we use two normal distributions for $S_{1}$ (black) and $S_{1}^{\text{COE}}$ (red). (a) Ratio continuously decreases. (b) Ratio stays one. (c) Ratio continuously increases.
Approach of the distributions? – When two distributions approach each other, this implies also an approach of their moments. Figure 2(a) shows the absolute distance for the first moment, i.e. the average of the von Neumann entropy for the KFIM and the COE (KumPan2011, , Eq. (48)) (black crosses), i.e.
$\Delta_{\overline{S_{1}}}=\left|\overline{S_{1}}-\overline{S_{1}^{\text{COE}}}\right|.$ (8)
In dependence on the chain length $L$ a decrease of $\Delta_{\overline{S_{1}}}$ is found. For $L>10$ it is well described by an exponential decay, as seen by the straight line behavior in the displayed semi-logarithmic representation. This clearly demonstrates, that the average eigenstate entanglement for the KFIM indeed approaches the COE prediction (KumPan2011, , Eq. (48)) in the limit of large dimensions.
In Figure 2(b) we see a similar behavior for the absolute distance of the second moment, i.e. the standard deviation of the distributions for KFIM and COE (black crosses), i.e.
$\Delta_{{\sigma}}=\left|\sigma(S_{1})-\sigma(S_{1}^{\text{COE}})\right|\;.$ (9)
Thus, also the standard deviation of the eigenstate entanglement of the KFIM approaches the COE result.
However, even if both average and standard deviation of two distributions approach each other, it is still possible, that the distribution remain distinct and do not approach each other:
Assuming two distributions with approaching average values and standard deviations three different situations can arise, as illustrated in Fig. 3. Figure 3(a) shows the case, for which the distributions actually approach each other. In (b) the distributions remain close but do not approach each other and in (c) the two distributions form distinct peaks. To distinguish between these three cases we use the ratio of the absolute distance between the average values to the standard deviation PauCarRodBuc2021
$\mathcal{R}=\Delta_{\overline{S_{1}}}/\sigma(S_{1})\;.$ (10)
In case (a) the two distributions approach each other. Thus this ratio is decreasing, i.e. the average values approach each other faster, than the decrease of the standard deviation. In case (b) the ratio (10) is always one, so the distributions are always close but there is no approach. In case (c) the ratio (10) is increasing, i.e. the widths of the distributions become smaller more quickly than the averages approach each other so that the distributions show distinct peaks.
We now apply this criterion to the KFIM. Figure. 2(c) shows the ratio (10) in dependence of the chain length $L$. Initially, the ratio (10) increases, while for large $L$ it appears to saturate with a value $\mathcal{R}\approx 1.25$ for $L=20$. This leads to the remarkable conclusion, that the distribution of the von Neumann entropy for the KFIM and the COE remain close but do not approach each other with increasing system size. This result is also supported by a similar behavior of the Kullback-Leibler-divergence KulLei1951 ; CovTho2006 and the Jensen-Shannon-divergence Lin1991 (not shown).
Therefore, the question for the reason of these deviations from the COE behavior arises. A first guess is that these are due to the local coupling applied in the KFIM. To test this, we discuss in the following two other kicked spin-$1/2$ chain models and a random matrix model for kicked many-body systems. One could expect, that with more complex coupling the deviations of the entropy from the COE result will decrease.
Further models – As second model we consider a slightly more complex coupling than the KFIM. The $C_{\text{NN}}$ chain is given by the time evolution operator SieLewScaZak2023
$U_{\text{$C_{\text{NN}}$}}=$ $\text{e}^{-\text{i}H^{x}_{\text{$C_{\text{NN}}$}}}\text{e}^{-\text{i}H^{z}}\quad\text{with}$ (11)
$H^{x}_{\text{$C_{\text{NN}}$}}=$ $\frac{g}{2}\sum_{j=1}^{L}\left(\sigma^{x}_{j}+\sigma^{x}_{j}\sigma^{x}_{j+1}\right)$ (12)
and $H^{z}$ as for the KFIM, see Eq. (6). Compared to the KFIM, the $C_{\text{NN}}$ chain has an additional nearest neighbor coupling in the $x$ component. The third model is the $C_{\text{NNNN}}$ chain SieLewScaZak2023
$U_{\text{$C_{\text{NNNN}}$ }}=$ $\text{e}^{-\text{i}H^{x}_{\text{$C_{\text{NNNN}}$ }}}\text{e}^{-\text{i}H^{z}}\quad\text{with}$ (13)
$H^{x}_{\text{$C_{\text{NNNN}}$ }}=$ $\frac{g}{2}\sum_{j=1}^{L}\left(\sigma^{x}_{j}+\sigma^{x}_{j}\sigma^{x}_{j+1}+\frac{2}{3}\sigma^{x}_{j}\sigma^{x}_{j+3}\right)$ (14)
and $H^{z}$ as for the KFIM, see Eq. (6). This model has an additional next–next–nearest neighbor coupling in the $x$ component. Additionally, we study a random matrix model for quantum chaotic Floquet many body systems given by FriKie2024 ,
$U_{\text{random matrix model}}=$ $U_{c}\left({U_{1}\otimes U_{2}\otimes\dots\otimes U_{L}}\right)$ (15)
$\left[U_{c}\right]_{kl}=$ $\exp\left(\text{i}\xi_{k}\right)\delta_{kl}\;,$ (16)
with $U_{j}$ for $j\in\{1,2,\dots,L\}$ being independent $2\times 2$ matrices chosen randomly from the CUE and $\xi_{k}$ for $k\in\{1,\dots,2^{L}\}$ being random variables uniformly distributed in $\left[-\pi,\pi\right[$. Similar to the spin-$1/2$ chain models, this model has two-dimensional local Hilbert spaces, but with a diagonal global coupling. Numerical paramters are given in Note1 .
Figure 2(a) shows for each model the absolute distance between the average von Neumann entropy and the COE prediction (8). All models show the same behavior, as already found for the KFIM: Namely, for all quantum chaotic spin-$1/2$ chains and the random matrix model the average eigenstate entanglement approaches the COE prediction (KumPan2011, , Eq. (48)). Figure 2(b) demonstrates that also the standard deviation of the von Neumann entropy approaches the COE result for all discussed models. Figure 2(c) shows the ratio (10) for all models. In neither case the ratio becomes small when increasing $L$, but instead approaches some constant between one and two. Thus for all considered quantum chaotic kicked spin-$1/2$ chains as well as for the random matrix model, the distribution of the eigenstate entanglement remains close to the distribution for the COE, but the two distributions do not approach each other.
Since this behavior occurs for all models with very different coupling ranges, the origin for the deviations cannot be the local coupling. Instead, the cause must be a property common to all models. Therefore we conclude that the deviations are due to the local two-dimensional Hilbert spaces of the models. The type of coupling is responsible for differences in the details of the behavior of ratio (10) and its approach towards saturation.
Discussion and outlook – We demonstrate that for prototypical examples of quantum chaotic kicked spin-$1/2$ chains the average eigenstate entanglement approaches the COE result when increasing the system size. However, we find that the distribution of the entanglement entropy remains significantly different from the distribution for the COE. These deviations for the entropy are demonstrated for three different kicked spin-$1/2$ chain models and a random matrix model, which takes into account the local two-dimensional Hilbert spaces of the chains but contains a global coupling. Thus the observed deviations are due to the local substructure, while the coupling of the subsystems affect the form of the deviations.
We have also performed numerical computations of the linear entropy (closely related to the purity), for which similar deviations of the statistics from the results for the COE are found. Therefore one could think that also other eigenvector statistics might be affected. It turns out that the multifractal dimensions of eigenstates, see e.g. BaeHaqKha2019 ; PauCarBucRod2021 and references therein, for all considered models approach the COE results. This means, that entanglement statistics is more sensitive to deviations from RMT than the multifractal dimensions.
In the future it would be interesting to understand the influence of the dimensions of the subsystems. For example, for the random matrix model it is observed FriKie2024 that for large subsystems the spectral form factor approaches the CUE instead of the COE. Possible systems to study this would have a larger spin of the subsystems, e.g. spin-1 or spin-2 chains. Another important question is whether and if so to what extent the deviations of the eigenstate entanglement statistics from the COE influence the dynamical behavior of the system. Similar deviations of the entanglement of time-evolved wave-packets could mean that thermalization properties will be different from the RMT expectations.
We thank Felix Fritzsch, Masud Haque, Roland Ketzmerick, Maximilian Kieler, and Ivan Khaymovich for useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 497038782.